If a company mistakenly counts less items during a physical inventory than actually exist,


Question 13 

 If a company mistakenly counts less items during a physical inventory than actually exist, how will the error affect the cost of goods sold? 

  • understated 

  • overstated 

  • no change 

  • only inventory will be affected


12. Problems and Applications Q12 

The economy of Elmendyn contains 900 $1 bills. 

If people hold all money as currency, the quantity of money is _______  

If people hold all money as demand deposits and banks maintain 100 percent reserves, the quantity of money is _______ 

If people hold equal amounts of currency and demand deposits and banks maintain 100 percent reserves, the quantity of money is _______  

If people hold all money as demand deposits and banks maintain a reserve ratio of 12.5 percent, the quantity of money is _______  

If people hold equal amounts of currency and demand deposits and banks maintain a reserve ratio of 12.5 percent, the quantity of money is _______ 
Answer #1

Part 13

Option A

understated

when physical inventory is counted less, the cost of goods sold is understated and net income is overstated

Part 12

a If people hold all money as currency, the quantity of money is $900.

b If people hold all money as demand deposits and banks maintain 100 percent reserves, the quantity of money is $900.

c If people hold equal amounts of currency and demand deposits and banks maintain 100 percent reserves, the quantity of money is $900.

d. If people hold all money as demand deposits and banks maintain a reserve ratio of 12.5 percent, the quantity of money is $7200 (1/12.5%*900)

e. If people hold equal amounts of currency and demand deposits and banks maintain a reserve ratio of 12.5 percent, the quantity of money is $12800

Quantity of money in case of all money held as demand deposits = 7200

8*(7200-C)= D

Thus,

8*(7200-D)= D

57600-8D=D

9D = 57600

D = 6400

Substituting D = 6400

8*(7200-C)=6400

7200-C = 800

C = 6400

Quantity of money = 6400+6400 = 12800

> e. is incorrect; the number is $1,600. I hope this helps:)

stella luvs Sun, Dec 5, 2021 9:38 PM